Circular No. 8769

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 U6 (SPACEWATCH) An apparently asteroidal object discovered by the Spacewatch pro-

gram, and posted on the Minor Planet Center’s NEOCP webpage, has been reported to be marginally cometary in appearance by observers elsewhere: E. Christensen reports that his inspection of 30-s CCD frames taken by A. Gibbs at Mt. Lemmon on Oct. 28.2 UT suggests that the object’s images appear soft. J. Young (Table Mountain) writes that his images, obtained in excellent seeing conditions (despite smoke from nearby fires) and moonlight on Oct. 30.1, show a diffuse 3′′-4′′ coma with no hint of a tail.

2006 UT α_{2000} δ_{2000} Mag. Observer
Oct. 19.13426 23^{h}32^{m}59.^{s}24 +3^{o}15′56′′.5 19.8 Spacewatch

The available astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-U79.

T = 2008 June 5.408 TT  ω = 276.^{o}510
Ω = 180.182 }2000.0
q = 2.50156 AU  i = 84.964

COMET C/2006 U7 (GIBBS) Alex R. Gibbs reports his discovery of a comet on images taken with

the Mt. Lemmon Survey 1.5-m reflector, the object then reported as diffuse with diameter 7′′, possibly elongated in p.a. 230^{o}. Following posting on the NEOCP, J. Young notes a very diffuse 4′′-5′′ coma and no tail on images taken at Table Mountain on Oct. 30.4 UT, adding that the comet is in a faint, somewhat-dense star field. Follow-up observations at Mt. Lemmon by Christensen on Oct. 30.4 (four 90-s exposures in good seeing) reveal a diffuse 5′′ coma and a 15′′ tail in p.a. 260^{o}.

2006 UT α_{2000} δ_{2000} Mag. Observer
Oct. 28.29549 3^{h}31^{m}54.^{s}84 +17^{o}55′24′′.0 20.8 Gibbs

The available astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-U80.

T = 2008 Feb. 23.969 TT  ω = 68.^{o}853
Ω = 56.166 }2000.0
q = 3.52395 AU  i = 11.253
2006 October 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green